OpenStudy (omgthatgirl):

Find the first three output values of the fractal generating function f(z)=z^2-3+2i a. -3+2i, 2-10i, -99-38i b. 2-10i, -99-38i, 8354+7526i c. -3+2i, 2-10i, 8354+7526i d. -3+2i, -34-129i, 8354+7526i

OpenStudy (perl):

to solve this problem you need to be given a 'seed' value

OpenStudy (omgthatgirl):

can u help me

OpenStudy (perl):

lets assume that z = 0 is your seed value

OpenStudy (perl):

$$ \Large { f(z) = z^2 -3 + 2i\\ \\ \therefore \\ f(0) = 0^2-3 + 2i = -3+2i\\ f(-3+2i) = ( -3+2i)^2 -3 + 2i = ? } $$

OpenStudy (perl):

are you allowed to use a calculator?

OpenStudy (omgthatgirl):

yes